Consider the Schrödinger equation for a particle in the presence of the potential
of the form
1
Expert's answer
2021-09-13T12:17:17-0400

"\frac{h^2d^2\psi(x)}{2m{dx^2}}+V(x)\psi(x)= E\psi(x)"


and

"\psi(x) = Ax{e^\frac{-x^2}{c^2}}" having "E = 0"

Therefore; "\frac{h^2d^2\psi(x)}{2m{dx^2}}+V(x)\psi(x)= 0"